# Are Large Language Models Rational Investors?

### v1

@@ Abstract
Large Language Models (LLMs) are progressively being adopted in financial analysis to harness their extensive knowledge base for interpreting complex market data and trends. However, their application in the financial domain is challenged by intrinsic biases (i.e., risk-preference bias) and a superficial grasp of market intricacies, underscoring the need for a thorough assessment of their financial insight. This study introduces a novel framework, Financial Bias Indicators (FBI), to critically evaluate the financial rationality of LLMs, focusing on their ability to discern and navigate the subtleties of financial information and to identify any irrational biases that might skew market analysis.
Our research adopts an innovative methodology to measure financial rationality, integrating principles of behavioral finance to scrutinize the biases and decision-making patterns of LLMs. We conduct a comprehensive evaluation of 19 leading LLMs, considering factors such as model scale, training datasets, input strategies, etc. The findings reveal varying degrees of financial irrationality among the models, influenced by their design and training. Models trained specifically on financial datasets might exhibit greater irrationality, and it’s possible that even larger financial language models (FinLLMs) could display more biases than smaller, more generalized models. This outcomes provide profound insights into how these elements affect the financial rationality of LLMs, indicating that targeted training and structured input methods could improve model performance. This work enriches our understanding of LLMs’ strengths and weaknesses in financial applications, laying the groundwork for the development of more dependable and rational financial analysis tools.
@@
Are Large Language Models Rational Investors?
Yuhang Zhou††thanks: Email: yuhangzhou22@m.fudan.edu.cn Yuchen Ni Xiang Liu Jian Zhang Sen Liu Guangnan Ye††thanks: Corresponding Author. Email: yegn@fudan.edu.cn Hongfeng Chai Institute of Fintech, Fudan University School of Computer Science, Fudan University School of Electronics and Information Engineering, Tongji University Tandon School of Engineering, New York University DataGrand Inc.
@@ 1 Introduction
Recent advancements in LLMs, such as GPT-4 OpenAI (2023) and LLaMA-2 Touvron et al. (2023), have shown their prowess across a spectrum of natural language processing (NLP) tasks Zhao et al. (2023), extending into specialized domains such as finance Zhang et al. (2023), law Cui et al. (2023), and healthcare Wang et al. (2023). Despite their versatility, these models grapple with inherent biases Gallegos et al. (2023), Rutinowski et al. (2023), encompassing gender, race, and socioeconomic disparities, which could compromise their reliability and entail significant consequences Jeoung et al. (2023). Efforts to mitigate such biases have led to the development of benchmark datasets such as StereoSet Nadeem et al. (2020) for stereotype identification and OpinionGPT Haller et al. (2023) for generating bias-neutral content, employing techniques such as LoRA Hu et al. (2021) for model fine-tuning, especially within the social sciences. As we delve into the application of these advanced models in specific sectors, it becomes imperative to address these biases to ensure fair and unbiased outcomes.
@@
In the realm of financial LLMs (FinLLMs), research has predominantly concentrated on enhancing model performance through continued pre-training or fine-tuning, with evaluation metrics focused on NLP tasks and financial applications. BloombergGPT Wu et al. (2023), a pioneering FinLLM, was trained on a vast financial corpus using BLOOM Workshop et al. (2022), and assessed across a broad array of financial and general NLP tasks. The FinQA dataset Chen et al. (2021), comprising QA pairs derived from S&P 500 company reports, served as a notable example of domain-specific evaluation. Moreover, LLMs have been employed in financial analytics, acting as advisors by leveraging news Lopez-Lira and Tang (2023) and fundamental analyses Fatouros et al. (2024) for investment decisions across diverse securities Romanko et al. (2023). However, the efficacy of these models is contingent upon the models’ own rationality as participants in the market. A lack of rationality in large models could lead to misinterpretations and misapplications of market dynamics, adversely impacting not only the users of these models but also the broader economy. Figure 1 is an example of model irrationality, where the model gives inconsistent expectations for the same event from different subjects, resulting in different emotions.
Hence, it is imperative to examine the rationality of LLMs before their integration into financial advisory roles. Limited research has addressed the financial biases in pre-trained models, with methodologies such as probabilistic detection Chuang and Yang (2022) and consistency checks Yang et al. (2023) being explored. However, these approaches primarily focus on data processing and lack a comprehensive financial theory to substantiate their findings, rendering them inadequate for assessing LLM financial rationality. This underscores the critical need for a holistic framework to evaluate the rationality of LLMs in financial contexts.
@@
To work towards this goal, our study conducts an examination of the financial rationality of LLMs based on the theory of behavioral finance Barberis and Thaler (2003). We believe that the enduring theory of behavioral finance, grounded in psychology and finance, can provide a more comprehensive perspective to support research findings. In the current scenario where it is challenging to quantify behavioral finance indicators, we propose the Financial Bias Indicators (FBI) framework, which provides a comprehensive assessment of financial rationality within LLMs from multiple perspectives.
Our framework specifically delineates the financial rationality of LLMs into two dimensions: Belief and Risk Preference. By leveraging theories on human cognitive biases, we assess LLMs’ rationality by observing their reactions to phenomena, e.g., limited attention bias, anchoring effect, and loss aversion, thereby gauging their rationality. Additionally, we delve into the models’ reasoning processes in financial contexts to distinguish whether observed biases arise from inherent capabilities or a lack of robustness. This approach enables us to discern the nuanced behaviors of LLMs, revealing a spectrum of financial irrationality across models of varying sizes. Notably, our research uniquely focuses on the biases of models within China’s A-share market, a domain scarcely explored in existing literature, thereby filling a critical gap in the understanding of LLM behavior in specific financial environments.
Our findings indicate that both larger and smaller Foundation Models111Foundation Models is an alias for Large Language Models, which is used interchangeably in this article. manifest degrees of financial irrationality, with larger models showing subtler biases and smaller ones more pronounced irrational behaviors across different companies, industries, and scenarios. This irrationality, potentially deepened by continuous pre-training or incremental fine-tuning with financial datasets, poses a risk of market anomalies when such models are applied in real-world financial scenarios. Despite the potential for mitigation through instructional methods or reasoning approaches such as Chain of Thought (COT) Wei et al. (2022), the inherent biases in LLMs, underscore the need for further research aimed at enhancing the models’ robustness, fairness, and rationality to safeguard financial market stability and protect personal assets.
@@
We have found that LLMs often exhibit irrational phenomena in the field of financial quantification research. So we introduce the Financial Bias Indicators (FBI) framework, a novel approach for assessing the rationality of LLMs in finance. Grounded in behavioral finance, the FBI framework categorizes financial rationality into Belief Bias and Risk-preference Bias, offering a new perspective for evaluating LLM rationality in financial contexts. To our best knowledge, this is the first study on the rationality of LLMs in the financial domain, and also the first study to quantify behavioral finance indicators, which can help future researchers develop more reliable FinLLMs (section 3). We have published our code and data on GitHub222https://github.com/zhiqix/FBI.
The FBI framework conducts an extensive analysis of the top 19 Chinese and English oriented LLMs, examining the impact of model parameters, training data, and input formats on their financial rationality. This comprehensive review provides insights into the varying degrees of financial irrationality across different models, enhancing our understanding of their behavior in financial applications(section 4 and section 5).
Based on the FBI framework, we further investigate the root causes of financial irrationality in LLMs from various perspectives. Gaining a deeper understanding of these factors enables more effective evaluation and mitigation of financial irrationality in LLMs, contributing to the creation of more robust and dependable models. This supports the advancement of global financial development(section 6).
@@ 2 Background
Behavioral finance Barberis and Thaler (2003), an area at the confluence of psychology and economics, investigates how psychological influences and cognitive biases affect the financial behaviors of individuals and institutions and, by extension, financial markets. This field seeks to elucidate the underlying psychological factors that lead to various market phenomena, diverging from traditional finance theories that assume rational market participants. By dissecting individual and collective financial decision-making processes, behavioral finance aims to construct a realistic framework of market dynamics, incorporating cognitive errors and the limits of arbitrage. We adopt the classification proposed by Barberis and Thaler (2003), segmenting behavioral finance into Cognitive Bias and Limits to Arbitrage, to structure our exploration of these dimensions.
@@ 2.1 Cognitive Bias
Cognitive biases represent systematic departures from normative decision-making, influencing how investors form beliefs and assess risks. These biases are multifaceted, manifesting as erroneous beliefs or inconsistent risk preference. For instance, belief biases, such as attentional neglect, representativeness bias, anchoring effect, and overconfidence,skew investors’ expectations, while risk-preference biases, such as loss aversion and reference dependence, lead to irregularities in risk assessment and decision-making under uncertainty. A comprehensive taxonomy of these biases, alongside their definitions, is detailed in Appendix B.
@@ 2.2 Limits to Arbitrage
Contrary to the Efficient Markets Hypothesis (EMH), which posits that asset prices fully reflect all available information and that market participants behave rationally, behavioral finance identifies scenarios where irrationality prevails, leading to anomalies like market bubbles and systemic crises. These phenomena are often attributed to the collective impact of cognitive biases on investor expectations, which can cause significant deviations from asset fundamentals. This section delves into how such irrational behaviors challenge the EMH, introducing inefficiencies that arbitrage mechanisms cannot always correct.
@@ 3 FBI: A Framework for Assessing LLMs Financial Rationality
We propose the FBI framework illustrated in Figure 2, drawing upon behavioral finance principles and the integration of NLP technologies in the financial sector. This framework bifurcates into two segments to analyzing biases in LLMs: Belief Bias and Risk-preferences Bias.
@@ 3.1 Belief Bias
In the contemporary milieu, the perpetual influx of novel information necessitates ongoing revisions to our forecasts and anticipations, shaping our beliefs. Within the belief bias examination, this framework endeavors to authenticate four biases: Anchoring Effect, Limited Attention Bias, Representativeness Bias, and Overconfidence, employing a uniform dataset. It evaluates the capacity of various LLMs to sustain objective rationality in actual market dynamics.
To evaluate the rationality of LLMs in financial markets comprehensively, we analyze the models’ responses to newly emerging information. This involves categorizing messages into two main types: event news and interactions, especially considering the noise prevalent in online investor dialogues within the Chinese market.
We adopt a refined classification of financial events, inspired by Zhou Zhou et al. (2021), into four primary categories: Corporate Governance and Equity Changes (CGEC), Financial Reports and Earnings Expectations (FREE), Market Behavior and Announcements (MBA), and Negative Events and Risk Management (NERM). This leads to a set of specific event types , as elaborated in Appendix B.
From January 2023 to December 2023, we compiled a corpus of news pieces, with a focus on including a subset that contains emotionally charged articles. This inclusion aims to enrich the LLM training dataset with event news that embodies emotional dimensions. The selected subset comprises 9 articles with positive emotions, 9 with negative emotions, and 6 with mixed emotions, resulting in a total of articles earmarked for detailed analysis.
Furthermore, we curated a set of 10 emotionally neutral question-response interaction pairs , to study LLMs’ comprehension in a controlled setting.
For each selected piece of news and interaction , we abstracted specific numerical details to proportional figures to standardize the data across varying scales of company market capitalizations.
We then sampled 600 companies from the Chinese A-share market, excluding those marked for delisting (ST), distributed equally across three tiers based on market capitalization: top, middle, and bottom, with companies from each tier. This selection was anonymized by replacing company and industry names to test the belief bias in LLMs, with a detailed classification by the first-level industry provided in Appendix C.
@@ 3.1.1 Analysis Logic
Our analysis framework is designed to unravel various belief biases potentially embedded within LLMs, using a structured approach to evaluate the models’ reactions to financial news across varying contexts.
Anchoring Effects
To detect anchoring effects, we examine whether LLMs display differing interpretations of identical events when presented with varying company contexts, indicative of bias ingrained from historical data. This involves modifying the subject company in each news item and assessing the variability in LLM evaluations using Analysis of Variance (ANOVA) St et al. (1989), denoted as , where tests the variance in scores across companies for a given news item . The specific prompt structure for LLMs is detailed in Appendix F.
Representativeness Bias
We explore the representativeness bias by correlating LLM output scores with company size and industry sector. The size correlation is quantified using the Spearman correlation coefficient , where represents the LLM scores and the market capitalizations of companies . Industry correlation analysis utilizes ANOVA, expressed as , alongside box plots to compare score distributions across different industries , representing the first-level industry classification of companies.
Overconfidence
To assess overconfidence, we observe score volatility for identical events with varied subjects among models undergoing continuous pre-training or fine-tuning on financial datasets. This is measured by the standard deviation of scores for the same event applied to different company contexts.
Limited Attention Bias
For detecting limited attention bias, we adopt a Comparative Output Testing (COT)-like methodology, allowing models to engage in a deliberative thought process before issuing scores. This process aims to discern if a more reflective approach versus a rapid response leads to more consistent scoring, analyzed through a variance comparison .
Score Instability Analysis
To investigate the roots of score instability, be it due to inadequate reasoning or compromised rationality, we utilize Bertopic Grootendorst (2022) for topic word extraction within the reasoning texts provided by the models. By clustering reasoning texts and extracting keywords, we analyze score discrepancies across different clusters, denoted by , to identify if certain thematic focuses lead to inconsistent evaluations.

### v2

# Are LLMs Rational Investors? A Study on Detecting and Reducing the Financial Bias in LLMs

@@ Abstract
Large Language Models (LLMs) are increasingly adopted in financial analysis for interpreting complex market data and trends. However, their use is challenged by intrinsic biases (e.g., risk-preference bias) and a superficial understanding of market intricacies, necessitating a thorough assessment of their financial insight. To address these issues, we introduce Financial Bias Indicators (FBI), a framework with components like Bias Unveiler, Bias Detective, Bias Tracker, and Bias Antidote to identify, detect, analyze, and eliminate irrational biases in LLMs. By combining behavioral finance principles with bias examination, we evaluate 23 leading LLMs and propose a de-biasing method based on financial causal knowledge. Results show varying degrees of financial irrationality among models, influenced by their design and training. Models trained specifically on financial datasets may exhibit more irrationality, and even larger financial language models (FinLLMs) can show more bias than smaller, general models. We utilize four prompt-based methods incorporating causal debiasing, effectively reducing financial biases in these models. This work enhances the understanding of LLMs’ bias in financial applications, laying the foundation for developing more reliable and rational financial analysis tools.
@@
Are LLMs Rational Investors?
A Study on Detecting and Reducing the Financial Bias in LLMs
Yuhang Zhou††thanks: Email: yuhangzhou22@m.fudan.edu.cn1,2 Yuchen Ni3 Yunhui Gan1,2 Zhangyue Yin1 Xiang Liu4 Jian Zhang5 Sen Liu1,2 Xipeng Qiu1 Guangnan Ye††thanks: Corresponding Author. Email: yegn@fudan.edu.cn1,2 Hongfeng Chai1,2 1School of Computer Science, Fudan University 2Institute of Fintech, Fudan University 3School of Electronics and Information Engineering, Tongji University 4Tandon School of Engineering, New York University 5DataGrand Inc.
@@ 1 Introduction
Recent advancements in LLMs, such as GPT-4 OpenAI (2023) and LLaMA-2 Touvron et al. (2023), have shown their prowess across a spectrum of natural language processing (NLP) tasks Zhao et al. (2023), extending into specialized domains such as finance Zhang et al. (2023), law Cui et al. (2023), and healthcare Wang et al. (2023). Despite their versatility, these models grapple with inherent biases Gallegos et al. (2023), Rutinowski et al. (2023), encompassing gender, race, and socioeconomic disparities, which could compromise their reliability and entail significant consequences Jeoung et al. (2023). Efforts to mitigate such biases have led to the development of benchmark datasets such as StereoSet Nadeem et al. (2020) for stereotype identification, GenderCare Tang et al. for gender bias detection, OpinionGPT Haller et al. (2023) for generating bias-neutral content, employing techniques such as LoRA Hu et al. (2021) for model fine-tuning, especially within the social sciences. As we delve into the application of these advanced models in specific domains, it becomes imperative to address these biases to ensure fair and unbiased outcomes.
@@
However, in the realm of financial LLMs (FinLLMs), research has predominantly concentrated on enhancing model performance through continued pre-training or fine-tuning, with evaluation metrics focused on NLP tasks and financial applications. Moreover, LLMs have been employed in financial analytics, acting as advisors by leveraging news Lopez-Lira and Tang (2023) and fundamental analyses Fatouros et al. (2024) for investment decisions across diverse securities Romanko et al. (2023). However, the efficacy of these models is contingent upon the models’ own rationality as participants in the market. A lack of rationality in LLMs could lead to misinterpretations and misapplications of market dynamics, adversely impacting not only the users of these models but also the broader economy. Figure 1 is an example of model irrationality.
Hence, it is essential to assess the rationality of LLMs before incorporating them into financial advisory roles. Limited research has addressed the financial biases present in pre-trained models, exploring methods such as probabilistic detection Chuang and Yang (2022) and consistency checks Yang et al. (2023). However, these methods mainly focus on pre-trained embedding models like BERTDevlin et al. (2019), limiting their application to LLM bias detection. The absence of evaluation standards for financial biases hinders thorough oversight, compromising assessment realism and objectivity. Consequently, there is an immediate need for a holistic framework to gauge LLMs rationality in finance.
Meanwhile, the detection of financial rationality and biases in LLMs faces four main challenges:
Q1: How to define the financial rationality and biases of LLMs? A theoretical framework is required to support the detection of LLMs.
Q2: How to detect and reveal financial biases in LLMs? A method needs to be developed to quantify theoretical indicators and construct relevant datasets.
Q3: How to investigate the origins of financial biases in LLMs? It is necessary to research whether these financial biases stem from the model’s capabilities or its robustness.
Q4: How to mitigate financial biases in LLMs? Methods need to be found that mitigate biases without compromising the original capabilities of the model.
@@
To work towards this goal, our study conducts an examination of the financial rationality of LLMs based on the theory of behavioral finance Barberis and Thaler (2003). We believe that the enduring theory of behavioral finance, grounded in psychology and finance, can provide a more comprehensive perspective to support research findings. In the current scenario where it is challenging to quantify behavioral finance, we propose the Financial Bias Indicators (FBI) framework to comprehensively assess financial rationality in LLMs. The FBI framework consists of four components: Bias Unveiler, Bias Detective, Bias Tracker, and Bias Antidote, covering the definition, detection, cause analysis, and mitigation of financial biases.
Our research shows that almost all LLMs exhibit financial irrationality. These biases, which may be exacerbated by continuous pre-training or fine-tuning with financial data, could lead to market anomalies in real-world applications. While prompt-based mitigation methods show promise, the persistent biases in LLMs highlight the necessity for further research to improve model robustness, fairness, and rationality, ensuring financial market stability and asset protection.
@@
The FBI framework, based on behavioral finance, defines, detects, analyzes, and mitigates financial biases in LLMs, offering a novel approach to evaluate their financial rationality. It is the first study to quantify behavioral finance indicators in LLMs, paving the way for more reliable LLMs.
The framework extensively analyzes 23 leading LLMs, assessing how model parameters, training data, and input formats affect financial rationality. Our study deepens the understanding of the varying levels of financial irrationality among models and their behavior in financial contexts.
Utilizing the FBI framework, we explore the origins of financial irrationality in LLMs, identify methods to mitigate bias, and develop a dataset of 200,000 financial causal texts named FinCausal to address biases with causal knowledge. Ultimately, we experimented with four prompt-based methods for bias mitigation, yielding encouraging results.
@@ 2 Background
Behavioral finance explores how psychological factors and cognitive biases influence the financial behaviors of individuals, institutions, and markets, differing from traditional finance theories that assume rational market participants. The field aims to uncover the psychological roots of various market phenomena, dissecting financial decision-making processes to build a realistic framework of market dynamics that includes cognitive errors and the constraints on arbitrage. We structure our investigation using the classification from Barberis and Thaler (2003), which divides behavioral finance into Cognitive Bias and Limits to Arbitrage.
@@ 2.1 Cognitive Bias
Cognitive biases represent systematic departures from normative decision-making, influencing how investors form beliefs and assess risks. These biases are multifaceted, manifesting as erroneous beliefs or inconsistent risk preference. For instance, belief biases, such as attentional neglect, representativeness bias, anchoring effect, and overconfidence,skew investors’ expectations, while risk-preference biases, such as loss aversion and reference dependence, lead to irregularities in risk assessment and decision-making under uncertainty. A comprehensive taxonomy of these biases, alongside their definitions, is detailed in Appendix C.
@@ 2.2 Limits to Arbitrage
Contrary to the Efficient Markets Hypothesis (EMH), which posits that asset prices fully reflect all available information and that market participants behave rationally, behavioral finance identifies scenarios where irrationality prevails, leading to anomalies like market bubbles and systemic crises. These phenomena are often attributed to the collective impact of cognitive biases on investor expectations, which can cause significant deviations from asset fundamentals.
@@ 3 FBI: A Framework for Assessing LLMs Financial Rationality
We propose the FBI framework illustrated in Figure 2. This framework is divided into four parts: Bias Unveiler defines financial biases in LLMs based on behavioral finance; Bias Detective constructs detection data and evaluates current leading LLMs for biases; Bias Tracker analyzes the causes of biases based on detection results and attention mechanisms; Bias Antidote builds a financial causal dataset and employs a series of prompt-based methods to mitigate bias phenomena.
4 Bias Unveiler
Based on the definitions from behavioral finance, we categorize biases in LLMs within financial contexts into Belief Bias and Risk-preference Bias. Within these categories, we define six related psychological biases.
@@ 4.1 Belief Bias
In today’s information-rich environment, constant updates require adjusting our predictions and beliefs. This framework investigates three cognitive biases—Anchoring, Representativeness, and Overconfidence—using real-world market data like news and shareholder discussions to test LLMs’ ability to maintain rationality in market conditions.
Anchoring Effects
We test LLMs for Anchoring Effects by checking if they show different views on the same event or give consistent responses under different company settings. This bias, derived from past data, can introduce sentiment analysis biases and potentially disrupt markets when used in finance.
Representativeness Bias
We investigate Representativeness Bias in LLMs by analyzing their outputs in relation to company size and sector. This bias towards size and industry can concentrate investment risks and cause problems.
Overconfidence
To measure overconfidence, we track score fluctuations for the same events with different subjects in FinLLMs and corresponding base LLMs. Aggressive scores with high deviation suggest overconfidence in these models’ event assessments or responses.
4.2 Risk-preference Bias
Asset returns are uncertain, affecting investor behavior based on risk-return preferences. Our study of Risk-preference Bias explains Situational Dependence Bias, loss aversion, and framing effect in various decision contexts, assessing LLMs’ risk preferences in different scenarios.
Situational Dependence Bias
Recognizing decision-making as a process shaped by previous experiences and contextual factors, we delve into the Situational Dependence Bias by examining if LLMs exhibit variable risk preferences across different scenarios.
Loss Aversion
Frame Effect
We investigate the Framing Effect by restating scenarios in various languages or expressions to track changes in LLM preferences, aiming to determine if linguistic framing influences LLM outputs, indicating bias in option presentation.
5 Bias Detective
5.1 Belief Bias
5.1.1 Data Desgin
To comprehensively assess the rationality of LLMs in financial markets, we scrutinized their responses to emergent information, distinguishing between event news and interactions, amid the prevalent noise in online investor dialogues in China. We analyzed historical events impacting company stock prices and adopted a refined classification of financial events into four primary categories: Corporate Governance and Equity Changes (CGEC), Financial Reports and Earnings Expectations (FREE), Market Behavior and Announcements (MBA), and Negative Events and Risk Management (NERM), detailed in Appendix C.
Throughout 2023, we compiled 300 news articles, including a subset of 24 emotionally nuanced pieces to enhance the bias detection in LLMs. This subset, detailed in Appendix C, contained articles categorized into nine positive, nine negative, and six mixed emotions.
Additionally, we assembled 10 neutral interaction pairs to analyze LLM comprehension in a controlled environment.
For each and , numerical details were abstracted to proportional figures to standardize data across varying company market caps, as documented in Appendix F.
We sampled 600 companies from the Chinese A-share market, excluding delisting entities, distributed across three tiers of market capitalization: top, middle, and bottom, each containing 200 companies. This selection method aimed to test Belief Bias in LLMs, with classifications by industry outlined in Appendix B.
@@ 5.1.2 Analysis Logic
To investigate Anchoring Effects, we altered the subject company in each news item and assessed variability in LLM evaluations using Analysis of Variance (ANOVA) St et al. (1989), expressed as , examining the variance in scores across companies for each news . Representativeness Bias was analyzed by correlating LLM outputs with company size and industry sector, using Spearman correlation coefficient , where and represent LLM scores and market capitalizations of companies , respectively. ANOVA was employed for industry correlation , with box plots to display score distributions across industries , reflecting first-level industry classifications.To measure Overconfidence, we used the standard deviation of scores , representing the variance in LLM evaluations of the same event across different company contexts, comparing results from FinLLMs with base LLMs.
5.1.3 Result
The evaluation of Belief Bias is principally conducted through the examination of event news and interactions. This analysis reveals a widespread Anchoring Effect across the majority of LLMs when the subjects of events and interactions are modified, with slight variations observed among different models. The average variance index, detailed in Figure 3, sheds light on the rationality levels of LLMs with respect to representativeness bias. Specifically, LLMs with a focus on the Chinese language, such as the GLM and Qwen family, exhibit commendable financial rationality, whereas the Xuanyuan and Baichuan family are more susceptible to irrational behavior.
In terms of Overconfidence, violin plots, presented in Appendix H.1, illustrate the score distributions of various texts across all models. A notable disparity is observed in the models’ responses to composite texts of positive and negative emotional content. As per Table 5, the GPT and InternLM models display a marked optimism, in contrast to the pronounced pessimism of the Qwen and GLM family. Furthermore, Figure 4 highlights that models trained on financial corpora experience a heightened score variability compared to their base counterparts.
In terms of Representativeness Bias, all LLMs exhibited a correlation coefficient below 10% between model scores and market capitalization, indicating a weak correlation. However, certain LLMs showed clear biases towards specific industries, as documented in Appendix H.1. For example, the FinQwen model consistently allocated lower scores to the Steel and Banking sectors. A comprehensive analysis reveals that the Media, Steel, Banking, and Non-Banking Finance sectors frequently occupy the extreme ends of the scoring spectrum across different models, whereas the Computer Science and Automobile sectors generally maintain a middle ground, exhibiting relative stability.